# local non-bossiness

### v4

@@ 6. Concluding remarks
In one-to-one matching problems, local non-bossiness only requires that student without school cannot modify the set of unassigned students without getting a place elsewhere. Thus, weak non-bossiness is stronger than local non-bossiness in this context.
Although any group strategy.proof mechanism is weak non-bossy, it follows from the proof of Lemma 3 that the combination of local non-bossiness and strategy-proofness does not implies on weak non-bossiness.
@@
A locally group strategy-proof mechanism that is locally bossy and weak bossy.
@@
We have that $\Omega(P)=((1,s_{1}),(2,s_{1}),(3,s_{2}))$ and $\Omega_{s_{1}}(P)=\{1,2\}$. If the student $1$ changes her preference to $P^{\prime}_{1}$ such that $P^{\prime}_{1}:s_{1},s_{2},s_{0}$, she remains assigned to $s_{1}$ but is left without schoolmates because $\Omega_{s_{1}}(P^{\prime}_{1},P_{2},P_{3})=\{1\}$. This argument also shows that $\Omega$ is weak bossy, because $\Omega_{s_{0}}(P)=\emptyset$ and $\Omega_{s_{0}}(P^{\prime}_{1},P_{2},P_{3})=\{2,3\}$. $\Box$
Although the school-optimal stable mechanism is not strategy-proof, it is non-bossy (Afacan and Dur, 2017). Hence, the 4 implies that local non-bossiness and strategy-proofness are independent properties.
The last examples show that local group non-bossiness is stronger than local non-bossiness (5), and than local group strategy-proofness is stronger that strategy-proofness (6).
@@
Denote by $top(P_{i})$ the most preferred alternative of student $i$ when her preferences are $P_{i}\in{\mathcal{L}}$. Let $\Omega:{\mathcal{P}}\rightarrow{\mathcal{M}}$ be the mechanism such that:
$\Omega(P)=\left\{\begin{array}[]{ll}((1,s_{1}),(2,s_{1}),(3,s_{2})),&\quad\quad\mbox{when}\,\,top(P_{1})=top(P_{2})=s_{1};\\ ((1,s_{1}),(2,s_{1}),(3,s_{1})),&\quad\quad\mbox{when}\,\,top(P_{1})=top(P_{2})=s_{2};\\ ((1,s_{0}),(2,s_{0}),(3,s_{0})),&\quad\quad\mbox{in another case}.\end{array}\right.$
It is not difficult to verify that $\Omega$ is locally non-bossy: the only two students who might change their partners without changing schools are $1$ and $2$. However, if only one of them changes her preferences, she will be assigned to $s_{0}$ (if she was originally assigned to $s_{1}$ or $s_{2}$), or will not change anyone’s assigned school otherwise.
@@
Denote by $top_{k}(P_{i})$ the $k$-th most preferred alternative of student $i$ under $P_{i}\in{\mathcal{L}}$. Let $\Omega:{\mathcal{P}}\rightarrow{\mathcal{M}}$ be the mechanism such that:
$\Omega(P)=\left\{\begin{array}[]{ll}((1,top_{1}(P_{1})),(2,s_{1}),(3,s_{2})),&\quad\quad\mbox{when}\,\,top_{2}(P_{1})=s_{2};\\ ((1,top_{1}(P_{1})),(2,s_{2}),(3,s_{1})),&\quad\quad\mbox{in another case}.\end{array}\right.$
It is easy to verify that $\Omega$ is strategy-proof. We claim that $\Omega$ is not locally group strategy-proof. Let $P=(P_{1},P_{2},P_{3})$ be a preference profile such that $P_{1}:s_{1},s_{2},s_{0}$ and $P_{2}:s_{2},s_{1},s_{0}$. Let $P^{\prime}_{1},P^{\prime}_{2}\in{\mathcal{L}}$ be such that $P^{\prime}_{1}:s_{1},s_{0},s_{2}$ and $P^{\prime}_{2}=P_{2}$. It follows that the coalition of schoolmates $\{1,2\}\in\Omega_{s_{1}}(P)$ can manipulate the mechanism $\Omega$ at $P$, because $\Omega_{1}(P)=s_{1}=\Omega_{1}(P^{\prime}_{1},P^{\prime}_{2},P_{3})$ and $\Omega_{2}(P^{\prime}_{1},P^{\prime}_{2},P_{3})=s_{2}P_{2}s_{1}=\Omega_{2}(P)$. $\Box$

### v5

@@ 6. Concluding remarks
In one-to-one matching problems, local non-bossiness only requires that a student without a school cannot modify the set of unassigned students without getting a place elsewhere. Thus, weak non-bossiness is stronger than local non-bossiness in this context. Furthermore, although every group strategy-proof mechanism is weakly non-bossy, it follows from the proof of Lemma 3 that the combination of local non-bossiness and strategy-proofness does not imply weak non-bossiness.
@@
A locally group strategy-proof mechanism that is locally bossy and weakly bossy.
@@
We have that $\Omega(P)=((1,s_{1}),(2,s_{1}),(3,s_{2}))$ and $\Omega_{s_{1}}(P)=\{1,2\}$. If student $1$ changes her preference to $P^{\prime}_{1}$ such that $P^{\prime}_{1}:s_{1},s_{2},s_{0}$, she remains assigned to $s_{1}$ but is left without schoolmates because $\Omega_{s_{1}}(P^{\prime}_{1},P_{2},P_{3})=\{1\}$. This argument also shows that $\Omega$ is weakly bossy, because $\Omega_{s_{0}}(P)=\emptyset$ and $\Omega_{s_{0}}(P^{\prime}_{1},P_{2},P_{3})=\{2,3\}$. $\Box$
Although the school-optimal stable mechanism is not strategy-proof, Afacan and Dur (2017) show that it is non-bossy. Hence, 4 implies that local non-bossiness and strategy-proofness are independent properties.
The next two examples show that local group non-bossiness is stronger than local non-bossiness, and that local group strategy-proofness is stronger that strategy-proofness.
@@
Denote by $top(P_{i})$ the most preferred alternative of a student $i$ when her preferences are $P_{i}\in{\mathcal{L}}$. Let $\Omega:{\mathcal{P}}\rightarrow{\mathcal{M}}$ be the mechanism such that:
$\Omega(P)=\left\{\begin{array}[]{ll}((1,s_{1}),(2,s_{1}),(3,s_{2})),&\quad\quad\mbox{when}\,\,top(P_{1})=top(P_{2})=s_{1};\\ ((1,s_{1}),(2,s_{1}),(3,s_{1})),&\quad\quad\mbox{when}\,\,top(P_{1})=top(P_{2})=s_{2};\\ ((1,s_{0}),(2,s_{0}),(3,s_{0})),&\quad\quad\mbox{otherwise}.\end{array}\right$.
It is not difficult to verify that $\Omega$ is locally non-bossy. Indeed, 11 and $2$ are the only two students who might change their schoolmates without changing schools. However, the definition of $\Omega$ ensures that, if $1$ or $2$ maintains her assignment after changing her preferences, then the whole matching does not change.
@@
Denote by $top_{k}(P_{i})$ the $k$-th most preferred alternative of a student $i$ under $P_{i}\in{\mathcal{L}}$. Let $\Omega:{\mathcal{P}}\rightarrow{\mathcal{M}}$ be the mechanism such that:
$\Omega(P)=\left\{\begin{array}[]{ll}((1,top_{1}(P_{1})),(2,s_{1}),(3,s_{2})),&\quad\quad\mbox{when}\,\,top_{2}(P_{1})=s_{2};\\ ((1,top_{1}(P_{1})),(2,s_{2}),(3,s_{1})),&\quad\quad\mbox{otherwise}.\end{array}\right$.
It is easy to verify that $\Omega$ is strategy-proof. We claim that $\Omega$ is not locally group strategy-proof. Let $P=(P_{1},P_{2},P_{3})$ be a preference profile such that $P_{1}:s_{1},s_{2},s_{0}$, $P_{2}:s_{2},s_{1},s_{0}$, and $P_{3}:s_{1},s_{2},s_{0}$. Let $P^{\prime}_{1},P^{\prime}_{2}\in{\mathcal{L}}$ be such that $P^{\prime}_{1}:s_{1},s_{0},s_{2}$ and $P^{\prime}_{2}=P_{2}$. It follows that the coalition of schoolmates $\{1,2\}\in\Omega_{s_{1}}(P)$ can manipulate the mechanism $\Omega$ at $P$, because $\Omega_{1}(P)=s_{1}=\Omega_{1}(P^{\prime}_{1},P^{\prime}_{2},P_{3})$ and $\Omega_{2}(P^{\prime}_{1},P^{\prime}_{2},P_{3})=s_{2}P_{2}s_{1}=\Omega_{2}(P)$. $\Box$
Appendix B On local non-bossiness under non-responsive preferences
Given a set $N$ of students and a set $S$ of schools, consider a context in which every school $s$ has a complete, transitive, and strict preference relation $P_{s}$ defined on $2^{N}$ (the collection of subsets of $N$). Let $C^{s}:2^{N}\rightarrow 2^{N}$ be the mapping that associates to each $N^{\prime}\subseteq N$ the most preferred subset of $N^{\prime}$ under $P_{s}$. We refer to $C^{s}$ as the choice function induced by $P_{s}$.
In this context, when students have preferences $(P_{i})_{i\in N}\in{\mathcal{P}}$, a matching $\mu$ is stable as long as the following properties hold:
    Individual rationality. For all $i\in N$ and $s\in S$, $\mu(i)R_{i}s_{0}$ and $C^{s}(\mu(s))=\mu(s)$.
    No blocking. There is no $(i,s)\in N\times S$ such that $sP_{i}\mu(i)$ and $i\in C^{s}(\mu(s)\cup\{i\})$.
Notice that, throughout the paper, we implicitly assume that schools have responsive preferences (Roth and Sotomayor, 1990), as the relevant information about how schools choose students from a pool of applicants is given by priority orders $(\succ_{s})_{s\in S}$ and capacities $(q_{s})_{s\in S}$. More precisely, when schools’ preferences are responsive, for each $s\in S$ and $N^{\prime}\in 2^{N}$, $C^{s}(N^{\prime})$ is composed of the $\min\{|N^{\prime}|,q_{s}\}$ students in $N^{\prime}$ with the highest priority under $\succ_{s}$.
Without responsiveness, the existence of a stable matching can be compromised. To avoid this possibility and to ensure that some classical properties of $\rm{DA}$} continue to hold, [^19] The mechanism $\rm{DA}$} can be naturally adapted to more general contexts by assuming that each school uses its choice function to evaluate proposals when the student-proposing deferred acceptance algorithm is implemented. it is usually assumed that every $P_{s}$ satisfies the following conditions:
  • Substitutability: Given $N^{\prime},N^{\prime\prime}\in 2^{N}$ with $N^{\prime}\subseteq N^{\prime\prime}$,
    $i\in N^{\prime},\,\,\,\quad i\in C^{s}(N^{\prime\prime})\,]\quad\quad\Longrightarrow\quad\quad i\in C^{s}(N^{\prime}).$
  • Law of aggregate demand: Given $N^{\prime},N^{\prime\prime}\in 2^{N}$ with $N^{\prime}\subseteq N^{\prime\prime}$, $|C^{s}(N^{\prime})|\leq|C^{s}(N^{\prime\prime})|.$
Under substitutability, if a student is chosen from a group of candidates, she will still be chosen from any subgroup of candidates. On the other hand, the law of aggregate demand ensures that the number of students chosen does not shrink when the pool of candidates expands. In every school choice context $N,S,(P_{s})_{s\in S}]$ in which $(P_{s})_{s\in S}$ are substitutable and satisfy the law of aggregate demand, $\rm{DA}$} is the student-optimal stable mechanism and is strategy-proof (cf., Roth, 1984a; Hatfield and Milgrom, 2005).
It is natural to ask whether $\rm{DA}$} is locally non-bossy in this more general framework. Unfortunately, as the following example illustrates, the local non-bossiness of $\rm{DA}$} cannot be ensured when we only know that schools’ preferences are substitutable and satisfy the law of aggregate demand, even when only one school has non-responsive preferences.